# Hyperparameter Importance Analysis for Multi-Objective AutoML

### v2

@@ 4.2 Experiment: Time
Figure 3: Results for the time experiment. The Pareto front is on the left (error vs. training time in seconds) with the red dot being the default performance. The MO-fANOVA results are in the middle and the MO-ablation path analysis is on the right. The x-axis corresponds to the weighting of the minimum error objective.
@@
The first experiment used the MNIST dataset with Multi-layer Perceptron (MLP) classifiers based on Sci-kit learn’s implementation [30]. The objective “training time” was measured by the wallclock time the classifier used for training. The error was measured by subtracting the accuracy on the test set from 1. The MLPs were configured with 50 maximum epochs and Adam as the DL optimizer. SMAC tuned the MLPs hyperparameter w.r.t. the configuration space in Table 2. Figure 3 shows the Pareto front (left), the MO-fANOVA (middle) and the MO-ablation path analysis (right) results for the time experiment. Note that for the x-axis only the weighting for the error objective is displayed. Since the total weighting always sums to 1, the opposite weighting applies to the respective other objective — in this case, training time (so $w_{2}$ = $1-w_{1}$). The ablation path analysis plot is stacked, with each segment representing the performance increase attributed to a specific hyperparameter for each weighting. The blue part shows the performance of the default configuration.
@@ 4.3 Experiment: Fairness
Figure 4: Results for the fairness experiment. The Pareto front is on the left (error vs. demographic parity) with the red dot being the default performance. The MO-fANOVA results is in the middle and the MO-ablation path analysis is on the right. The x-axis corresponds to the weighting of the minimum error objective.
The second experiment used the Adults Census Income dataset [2] which contains several variables about US citizens with the binary target variable annual income higher or lower than 50’000\$. We only used the numeric and binary nominal variables of the dataset. It contains the sensitive variables “sex” and “race”. We calculated the fairness loss as the second objective for the experiment based on the prediction of the “race” variable using demographic parity (DP). It is calculated by the absolute difference of the mean proportions of positive predictions $y$ in each group, where the groups are defined by a sensitive variable $s$.
$DP=\left|\frac{\sum_{i=1}^{n}y_{{}_{i}}(s_{i}=0)}{n}-\frac{\sum_{i=1}^{n}y_{{}_{i}}(s_{i}=1)}{n}\right|$ (2)
The MLPs were set up the same as in the time experiment. Figure 4 shows the Pareto front, the MO-fANOVA, and the MO-ablation path analysis results for the fairness experiment. While the optimization found several fair configurations, an error remained, with a top accuracy of around 83%.
@@ 4.4 Experiment: Energy Consumption
Our last experiment used the Torchvision implementation of ResNets [16] with the first and the last layer set to a size of 3. The models were trained with AdamW and a cosine annealing learning rate scheduler with 200 maximum iterations. They were trained for 50 epochs with early stopping which monitored the test accuracy with a patience of 6 epochs. The CIFAR10 dataset was always normalized with mean and standard deviation. Optionally, the data was augmented with a random crop of size 32 and a random horizontal flip. The objective “inference energy”, so for predicting the test set, was estimated with CodeCarbon [25, 27] [^1] https://github.com/mlco2/codecarbon. The configuration space is displayed in Table 3. Figure LABEL:fig:energy shows the Pareto front, the MO-fANOVA and the MO-ablation path analysis results for the energy experiment.
@@ 5 Discussion
Our research aimed to investigate whether surrogate-based HPI methods could be converted to meaningfully measure HPI within MOO. We introduced the concept of weighting the objectives based on configurations along the Pareto front and applied it to two HPI methods: fANOVA and ablation path analysis. Subsequently, we conducted three MO HPO experiments with diverse datasets, models, and objectives to assess the usefulness of the two MO-HPI approaches. In the following, we discuss our research questions related to the experimental results from the last section.
@@ 5.1 Are the results intuitively correct?
The results show the effectiveness of our method. All three experiments provide plots that make sense from the ML perspective, and thus, we conjecture that the results are intuitively correct. This can be seen in Figure 3 for the ablation and LABEL:fig:energy for fANOVA, where hyperparameters related to network size (such as layer size and number of layers) strongly influence training time and energy consumption, consistent with the understanding that larger neural networks require more energy and time. A similar observation can be made for the learning rate. Its impact on the error is to be expected, given its well-known importance for achieving high accuracy. Additionally, in the second experiment, the learning rate plays a role in fairness (cf. Figure 4). The same is valid for data augmentation in the energy experiment, as is shown in Figure LABEL:fig:energy, it has a high importance and it is known to heavily influence performance [37]. However, some unexpected results, like the high importance of beta_1 in the fairness experiment, are not intuitively explainable and warrant further exploration. We note that recent results showed that improved fairness can be strongly related to hyperparameters [11]. Overall, we therefore claim that the results are intuitively correct, which positively answers our first research question.
@@ 5.2 How do the proposed methods compare?
When comparing the two methods, i.e. MO-fANOVA and MO-ablation path analysis, some interesting observations can be made. In the time experiment, the activation function is less important when measured by the ablation path analysis than by fANOVA, but epsilon gains more importance for the error objective in the ablation path analysis. Another surprising observation is that in the fairness experiment, beta_1 is deemed important by both measures but for the fANOVA measure it has a high influence on the error side, and for the ablation path analysis, it has a higher influence on the fairness. Another difference is the high importance of the learning rate in Figure 4 for fANOVA, while there is not much contribution of the learning rate visible for the ablation path analysis. In the energy experiment, in the ablation path analysis data augmentation rated high and the learning rate has minimal relevance in comparison to fANOVA.
@@
The difference in the results of the methods is mostly due to their distinct nature. It is well known that the ablation paths struggle with correctly attributing importance in case of interactions of hyperparameters [14]. Since the value of two or more hyperparameters has to be changed before an effect on the performance can be observed, the importance is only attributed to the last changed hyperparameter. Consequently, the performance differences upon changing the other hyperparameters might be lower than they actually are. In the single objective scenario, this is partially mitigated by also displaying the order of changes which is not possible with our visualisation.
@@
Furthermore, the ablation results strongly depend on the chosen default configuration as a reference point. If the default configuration is already fairly close to one of the extreme ends of the Pareto front, there are little performance gains that can be attributed to the hyperparameters. Some hyperparameters might even seem to be unimportant because they are already set to the optimal value with respect to the dominating objective. Therefore, it is important to always interpret the ablation path in view of the default configuration. Since fANOVA for main effects studies the variance for each hyperparameter independent of the others, it is more stable in that sense. Our method allows us to also study the importance of higher-order interaction effects with fANOVA, as in the single-objective case. To not clutter our plots and results, we have omitted this here.
@@
Lastly, it is important to note that ablation paths are considered a local HPI analysis since their path through the configuration space potentially only covers a small part of it. In contrast, fANOVA covers the entire configuration space and thus provides more high-level insights, which, on the other hand, might not be important for specific performance improvements w.r.t. the default configuration. Overall, it is helpful to use both methods, as they provide diverse insights.
@@ 5.3 Can we gain new insights that previous methods could not provide?
One could simply run HPI studies on each objective independently and compare their results. We argue that by considering different tradeoffs of the objectives – as typically done in MOO – we gain valuable insights with our methods that were not possible previously.
Considering the tradeoffs between the objectives, we can make the following observations. In the fANOVA plot of the time experiment (cf. Figure 3), most hyperparameters significantly impacting only one of the objectives intersect at approximately an equal weighting of the objectives. This suggests that the MO-fANOVA in this case contains as much information as calculating fANOVA independently. Nevertheless, even in Figure 3, it is not a trivial linear trend of the importance values of the hyperparameters over different objective tradeoffs. In Figure 4 for the fairness experiment, the intersections are still at around the same position but are not in the center of the plot. For the energy experiment (cf. Figure LABEL:fig:energy), the intersections are even at different x-positions.
In addition, it is evident that certain hyperparameters lose importance rapidly (e.g. the size of Layer 1 in Figure LABEL:fig:energy for the ablation path analysis), while others remain important for longer before they become irrelevant (e.g. beta_1 in Figure 4 for the ablation path analysis). Moreover, some hyperparameters would not be tuned when only looking at the objectives independently. For instance, in the ablation path analysis in Figure 4, the initial learning rate only contributes significantly to the performance within the plot but not on the edges.
Furthermore, it is interesting to observe that certain hyperparameters consistently retain importance, for instance, in Figure 4, for fANOVA both the initial learning rate and the number of neurons remain influential. Conversely, some hyperparameters remain unimportant across all weightings. This consistency provides clarity regarding whether these hyperparameters require tuning. Finally, the stacked visualization makes it possible to easily see the relative contribution of a hyperparameter for different points on the Pareto front which make up the weighting. Therefore, we conclude that there is more information in the plots that considers different objective tradeoffs compared to analyzing the objectives independently.
@@ 5.4 Limitations and Future Work
As always being a problem with post-hoc analysis, although our results are reasonable, the absence of ground truth data prevents us from guaranteeing their correctness. Nevertheless, we believe that there is sufficient empirical evidence allowing us to conjecture that our method is correct. Although several approaches for studying (single-objective) results of HPO are already presented [20, 4, 28, 38, 1, 32], a systematic set of criteria as proposed for interpretable ML [29] is still missing in the context of interpretable AutoML.
In our experiments, we have not explored MOO scenarios with more than two objectives. Although our approach can be easily extended to more objectives, plotting the results will be possible for three objectives in a 3D plot, but not beyond this. We note that many-objective optimization (i.e., more than four objectives) could be relevant in practice but leads to having a large fraction of all configurations being on the Pareto front. To the best of our knowledge, there are no reasonable solutions for many-objective HPO to date.
In future work, MO-AutoML could incorporate the proposed method to actively choose the most important hyperparameters to tune. This could be done in the MO setting by calculating the integrals of the MO-HPI over the different tradeoffs. This provides a way to quantify the importance one-dimensionally without having to decide on a tradeoff. This will lead to more efficient AutoML that learns to actively consider the HPI for different tasks, datasets, and models. Moreover, it is already known that post-hoc analysis with partial dependence plots of HPO runs can be skewed because HPO approaches, such as Bayesian optimization [34], tradeoff exploitation and exploration and thus bias the sampled configurations accordingly [28]. So far, it is not known how this affects approaches such as fANOVA and ablation paths, but it is reasonable to assume that they are similarly negatively affected. Therefore, future work has to include how this will affect our approach and new approaches for how to fix skewed results under biased data. In addition, our methods should be integrated into interactive tools, such as DeepCave [33] or IOHprofiler [10] where the weighting could be adjustable by a slider, and other parameters of the method could be adjustable. Another functionality could be enabling users to select configurations on the Pareto front and convert them into objective weightings.
@@ 6 Conclusion
In multi-objective hyperparameter optimization, assessing the importance of hyperparameters is challenging due to conflicting objectives. Our proposed method leverages surrogate-based hyperparameter importance measures, specifically fANOVA and ablation paths, in conjunction with a-priori scalarization across various objective tradeoffs. While we focus on those two HPI methods, our approach can also work with other HPI methods, such as investigating the interaction effects of hyperparameters. Empirical evaluations demonstrate the effectiveness of our method. We believe that our proposed approach will enhance the analysis of multi-objective AutoML results.
@@
The DFKI Niedersachsen (DFKI NI) is funded in the "zukunft.niedersachsen" by the Lower Saxony Ministry of Science and Culture and the Volkswagen Foundation (funding no. ZN3480). Marius Lindauer was supported by the German Federal Ministry of the Environment, Nature Conservation, Nuclear Safety and Consumer Protection (GreenAutoML4FAS project no. 67KI32007A).

### v3

@@ 4.2 Experiment: Time
The first experiment used the MNIST dataset with Multi-layer Perceptron (MLP) classifiers based on Sci-kit learn’s implementation [33]. The objective “training time” was measured by the wallclock time the classifier used for training. The error was measured by 1 minus the accuracy on the test set. The MLPs were configured with 50 maximum epochs and Adam as the DL optimizer. SMAC tuned the MLPs hyperparameter w.r.t. the configuration space in Table 2. Figure 3 shows the Pareto front (left), the MO-fANOVA (middle), and the MO-ablation path analysis (right) results for the time experiment. The Pareto front shows the set of non-dominated solutions. The red point represents the default configuration. Note that for the x-axis, only the weighting for the error objective is displayed. Since the total weighting always sums to 1, the opposite weighting applies to the respective other objective — in this case, training time, so $w_{2}$ = $1-w_{1}$. MO-fANOVA thus displays the HPI from a low weighting of the error, so a high weighting of the time objective, to a high weighting of the error and a low weighting of time. For example, the number of neurons starts at an importance of around 0.4 when only considering the time objective and decreases the more the error objective is weighed in. The ablation path analysis plot is stacked, with each segment representing the performance increase attributed to tuning a specific hyperparameter for each weighting. The blue part shows the performance of the default configuration. Note that the y-axis only starts at 0.9, which means the default configuration already does quite well without tuning the hyperparameters.
@@ 4.3 Experiment: Fairness
Figure 4: Results for the fairness experiment. The Pareto front is on the left (error vs. demographic parity loss), with the red dot being the default performance. The MO-fANOVA results are in the middle, and the MO-ablation path analysis is on the right. The x-axis corresponds to the weighting of the minimum error objective.
The second experiment used the Adults Census Income dataset [2], which contains several variables about US citizens with the binary target variable annual income higher or lower than 50’000\$. We only used the numeric and binary nominal variables of the dataset. It contains the sensitive variables “sex” and “race”. We calculated the fairness loss as the second objective for the experiment based on the “race” variable prediction using demographic parity (DP). It is calculated by the absolute difference of the mean proportions of positive predictions $y$ in each group, where the groups are defined by a sensitive variable $s$.
$DPLoss=\left|\frac{\sum_{i=1}^{n}y_{{}_{i}}(s_{i}=0)}{n}-\frac{\sum_{i=1}^{n}y_{{}_{i}}(s_{i}=1)}{n}\right|$ (2)
The MLPs were set up the same as in the time experiment. Figure 4 shows the Pareto front, the MO-fANOVA, and the MO-ablation path analysis results for the fairness experiment. While the optimization found several fair configurations, an error remained, with a top accuracy of around 83%. The plots can be interpreted the same way as described in section 4.2.
@@ 4.4 Experiment: Energy Consumption
Our last experiment used the Torchvision implementation of ResNets [17] with the first and the last layer set to a size of 3. The models were trained with AdamW and a cosine annealing learning rate scheduler with 200 maximum iterations. They were trained for 50 epochs with early stopping, which monitored the test accuracy with a patience of 6 epochs. The CIFAR10 dataset was always normalized with mean and standard deviation. Optionally, the data was augmented with a random crop of size 32 and a random horizontal flip. The objective “inference energy”, so for predicting the test set, was estimated with CodeCarbon [26, 29] [^1] https://github.com/mlco2/codecarbon. The configuration space is displayed in Table 3. Figure LABEL:fig:energy shows the Pareto front, the MO-fANOVA, and the MO-ablation path analysis results for the energy experiment. The plots can be interpreted the same way as described in section 4.2.
@@ 5 Discussion
Our research aimed to investigate whether surrogate-based HyperParameter Importance (HPI) methods could be converted to meaningfully measure HPI within MOO. We introduced the concept of weighting the objectives based on configurations along the Pareto front and applied it to two HPI methods: fANOVA and ablation path analysis. Subsequently, we conducted three MO-HPO experiments with diverse datasets, models, and objectives to assess the usefulness of the two MO-HPI approaches. In the following, we discuss our research questions related to the experimental results from the last section.
@@ 5.1 Are the results intuitively correct?
The results show the effectiveness of our method. All three experiments provide plots that make sense from the ML perspective, and thus, we conjecture that the results are intuitively correct. This can be seen in Figure 3 for the ablation and LABEL:fig:energy for fANOVA, where hyperparameters related to network size (such as layer size and number of layers) strongly influence training time and energy consumption, consistent with the understanding that larger neural networks require more energy and time. A similar observation can be made for the learning rate. Its impact on the error is to be expected, given its well-known importance for achieving high accuracy. Additionally, in the second experiment, the learning rate plays a role in fairness (cf. Figure 4). The same is valid for data augmentation in the energy experiment, as is shown in Figure LABEL:fig:energy, it has a high importance, and it is known to heavily influence performance [43]. However, some unexpected results, like the high importance of beta_1 in the fairness experiment, are not intuitively explainable and warrant further exploration. We note that recent results showed that improved fairness can be strongly related to hyperparameters [11]. Overall, we therefore claim that the results are intuitively correct, which answers our first research question positively.
@@ 5.2 How do the proposed methods compare?
When comparing the two methods, i.e., MO-fANOVA and MO-ablation path analysis, some interesting observations can be made. In the time experiment, the activation function is less important when measured by the ablation path analysis than by fANOVA, but epsilon gains more importance for the error objective in the ablation path analysis. Another surprising observation is that in the fairness experiment, beta_1 is deemed important by both measures, but for the fANOVA measure, it has a strong influence on the error side, and for the ablation path analysis, it has a stronger influence on the fairness. Another difference is the high importance of the learning rate in Figure 4 for fANOVA. At the same time, there is not much contribution of the learning rate visible for the ablation path analysis. In the energy experiment, in the ablation path analysis, data augmentation rated high, and the learning rate has minimal relevance in comparison to fANOVA.
@@
The difference in the results of the methods is mostly due to their distinct nature. It is well known that the ablation paths struggle with correctly attributing importance in case of interactions of hyperparameters [14]. Since the value of two or more hyperparameters has to be changed before an effect on the performance can be observed, the importance is only attributed to the last changed hyperparameter. Consequently, the performance differences upon changing the other hyperparameters might be lower than they actually are. In the single objective scenario, this is partially mitigated by also displaying the order of changes, which is not possible with our visualization.
@@
Furthermore, the ablation results strongly depend on the chosen default configuration as a reference point. If the default configuration is already fairly close to one of the extreme ends of the Pareto front, there are little performance gains that can be attributed to the hyperparameters. Some hyperparameters might even seem unimportant because they are already set to the optimal value concerning the dominating objective. Therefore, it is important to always interpret the ablation path in view of the default configuration. Since fANOVA for main effects studies the variance for each hyperparameter independent of the others, it is more stable in that sense. Our method also allows us to study the importance of higher-order interaction effects with fANOVA, as in the single-objective case. To avoid cluttering our plots and results, we have omitted this here.
@@
Lastly, it is important to note that ablation paths are considered a local HPI analysis since their path through the configuration space potentially only covers a small part of it and can be seen as a local interpretation of the incumbent configuration given a default configuration. In contrast, fANOVA covers the entire configuration space and thus provides more high-level insights, which, on the other hand, might not be important for specific performance improvements w.r.t. the default configuration. Overall, it is helpful to use both methods, as they provide diverse insights.
@@ 5.3 Can we gain new insights that previous methods could not provide?
We could simply run HPI studies on each objective independently and compare their results. We argue that by considering different tradeoffs of the objectives – as typically done in MOO – we gain valuable insights from our methods that were not possible previously.
Considering the tradeoffs between the objectives, we can make the following observations. In the fANOVA plot of the time experiment (cf. Figure 3), most hyperparameters significantly impacting only one of the objectives intersect at approximately an equal weighting of the objectives. This suggests that the MO-fANOVA in this case contains as much information as calculating fANOVA independently. Nevertheless, even in Figure 3, it is not a trivial linear trend of the importance values of the hyperparameters over different objective tradeoffs. In Figure 4 for the fairness experiment, the intersections are still at around the same position but are not in the center of the plot. For the energy experiment (cf. Figure LABEL:fig:energy), the intersections are even at different x-positions (i.e., objective weightings).
In addition, it is evident that certain hyperparameters lose importance rapidly (e.g., zero_init_residual in Figure LABEL:fig:energy for the ablation path analysis), while others remain important for longer before they become irrelevant (e.g., beta_1 in Figure 4 for the ablation path analysis). Moreover, some hyperparameters would not be tuned when only looking at the objectives independently. For instance, in the ablation path analysis in Figure 4 (right), the initial learning rate only contributes significantly to the performance for certain tradeoffs but not to one of the extrema of the Pareto front.
Furthermore, it is interesting to observe that certain hyperparameters consistently retain importance, for instance, in Figure 4, for fANOVA, both the initial learning rate and the number of neurons remain influential. Conversely, some hyperparameters remain unimportant across all weightings. This consistency provides clarity regarding whether these hyperparameters require tuning. Finally, the stacked visualization makes it possible to easily see the relative contribution of a hyperparameter for different points on the Pareto front, which make up the weighting. Therefore, we conclude that there is more information in the plots that consider different objective tradeoffs compared to analyzing the objectives independently.
@@ 5.4 Limitations and Future Work
As is always the problem with post-hoc analysis, although our results are reasonable, the absence of ground truth data prevents us from guaranteeing their correctness. Nevertheless, we believe that there is sufficient empirical evidence allowing us to conjecture that our method is correct. Although several approaches for studying (single-objective) results of HPO are already presented [21, 4, 30, 44, 1, 37], a systematic set of criteria as proposed for interpretable ML [32] is still missing in the context of interpretable AutoML.
In our experiments, we have not explored MOO scenarios with more than two objectives because visualization and interpretability would become complex. Nonetheless, our approach can be extended to more objectives. Consider that the weighting scheme is calculated by dividing each coordinate value of a point on an n-dimensional Pareto front with the total sum of all values of that point, with each objective previously normalized between 0 and 1. For both methods, the target Y can now be calculated as the weighted sum of objectives with which the surrogate models can be trained. Plotting the results for three objectives in a 3D plot would be possible, but not trivially beyond that. We note that many-objective optimization (i.e., more than three objectives) could be relevant in practice but lead to a large fraction of the configurations being on the Pareto front. To the best of our knowledge, there are no reasonable approaches for many-objective HPO to date.
In future work, MO-AutoML could incorporate the proposed method to actively choose the most important hyperparameters to tune. This could be done in the MO setting by calculating the integrals of the MO-HPI over the different tradeoffs. This provides a way to quantify the importance one-dimensionally without having to decide on a tradeoff. This will lead to more efficient AutoML that learns to actively consider the HPI for different tasks, datasets, and models. Moreover, it is already known that post-hoc analysis with partial dependence plots of HPO runs can be skewed because HPO approaches, such as Bayesian optimization [40], tradeoff exploitation and exploration and thus bias the sampled configurations accordingly [30]. So far, it is not known how this affects approaches such as fANOVA and ablation paths, but it is reasonable to assume that they are similarly negatively affected. Therefore, future work has to include how this will affect our approach and new approaches for how to fix skewed results under biased data. In addition, our methods should be integrated into interactive tools, such as DeepCave [38] or IOHprofiler [10]. Another functionality could be enabling users to select configurations on the Pareto front and convert them into objective weightings.
@@ 6 Conclusion
In multi-objective hyperparameter optimization, assessing the importance of hyperparameters is challenging due to conflicting objectives. Our proposed method leverages surrogate-based hyperparameter importance measures, specifically fANOVA and ablation paths, in conjunction with a-priori scalarization across various objective tradeoffs. While we focus on those two HPI methods, our approach can also work with other HPI methods, such as Partial Dependence Plots [30], symbolic regression [39], variants of fANOVA [44], or maybe even with visualization methods such as parallel coordinate plots [16] or configuration footprints [5, 36]. Empirical evaluations demonstrate the effectiveness and interpretability of our method. We believe our proposed approach will enhance the analysis of multi-objective AutoML results and thus contribute to a human-centered AutoML paradigm [28].
@@
The DFKI Niedersachsen (DFKI NI) is funded in the "zukunft.niedersachsen" by the Lower Saxony Ministry of Science and Culture and the Volkswagen Foundation (funding no. ZN3480). Marius Lindauer acknowledges support by the German Federal Ministry of the Environment, Nature Conservation, Nuclear Safety and Consumer Protection (GreenAutoML4FAS project no. 67KI32007A) and was funded by the European Union (ERC, “ixAutoML”, grant no.101041029). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. [Uncaptioned image]